Fermi National Laboratory
 

Accelerator Update for July 26, 2006 - July 28, 2006

 

Wednesday July 26

The day shift began with Operations monitoring store 4857 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, and with MiniBooNE taking beam.

Operations prepared the Recycler at 9:29 AM to begin stashing antiprotons. The stash ended and Pbar resumed stacking at 9:45.

At 10:13 AM, a Linac RF station (LRF5) tripped off. It reset okay.

Pbar experts began stacking studies at 10:52 AM.

At 12:27 PM, LRF5 tripped off for the first of two times in fifty-two minutes. It reset both times.

NuMI experts began turning on their power supplies at 2:43 PM.

The Recycler began preparing to stash at 2:44 PM. The stash ended and Pbar resumed stacking at 3:36 PM.

Operations established beam to NuMI at 4:44 PM. Experts began a study period.

The NuMI horn tripped off at 5:38 PM. It reset okay. It was not a LCW trip.

The Recycler began preparing to stash at 6:52 PM. The stash ended and Pbar resumed stacking at 7:13 PM.

At 7:37 PM, NuMI tripped off due to two quadrupole power supplies tripping off (QD101 & QD101F). Everything reset.

At 7:58 PM, the NuMI Lambertsons tripped off and would not reset. Operations called in a power supply expert.

The power supply expert reported at 8:16 PM that LAM61 had tripped on a Filter Rack Summation. After further investigation he reported that at least 30 capacitors were bad. At 9:15 PM, he reported that he would disconnect the filter cabinet so the experiment could run for the night. The bad capacitors will be replaced in the morning.

NuMI resumed taking beam at 9:27 PM.

At 9:44 PM, LAM61 caused the NuMI beam permit to trip, but the power supply did not trip. The permit reset.

The Recycler began preparing to stash at 11:29 PM.

At 11:59 PM, LAM60 tripped off. It reset okay.

Thursday July 27

The midnight shift began with Operations monitoring store 4857, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:52 AM, the NuMI beam permit tripped off due to beam in the kicker gap. It reset. At the same time a Pbar Debuncher septa (D:ISEP1V) tripped off. It also reset okay.

At 1:45 AM, NuMI tripped off due to excessive 60hz on LAM60. It reset.

Operations terminated store 4857 at 2:02 AM. Operations began preparing CDF and D0 for access into their collision halls.

At 2:29 AM, the NuMI beam permit tripped off for the first of four times in fifteen minutes on LAM60 excessive 60hz. It reset each time.

At 3:31 AM, NuMI tripped off due to QD101. It reset okay.

CDF came out of access at 3:47 AM.

At 3:52 AM, NuMI tripped off due to excessive 60hz on a correction magnet power supply (V100). It reset.

D0 came out of access at 4:39 AM.

At 4:43 AM, NuMI tripped off due to excessive 60hz on LAM60. It reset.

Operations put the Tevatron (TeV) through a dry squeeze at 5:21 AM.

Operations put the TeV into shot setup at 5:46 AM.

At 6:08 AM, NuMI tripped off for the first of two times in twenty-eight minutes due to QD101. It reset okay.

After resolving a closure problem, Operations established store 4859 at 8:11 AM with an initial luminosity of 171.2E30. This was the second best luminosity since shutdown.

The Neutron Therapy Facility (NTF) reported at 8:15 AM that they would be treating patients this afternoon.

A Controls expert reported at 9:03 AM that he had completed work on the NuMI bridge box.

The Recycler began preparing to stash at 9:07 AM. The stash ended at 9:43 AM.

At 10:59 AM, NuMI tripped off due to beam position problems. It reset okay.

At 12:29 PM, a MI power supply (South Anode) tripped off due to an interlock indication. Operations turned off the MI beam while an expert investigated. He fixed a failed resistor in MIRF3. Beam resumed at 12:57 PM.

NuMI turned off at 1:16 PM so experts could begin work on the LAM60 filters.

The Recycler began preparing to stash at 1:47 PM. The stash ended at 2:15 PM.

The LAM60 work was completed and NuMI resumed taking beam at 3:07 PM.

At 4:18 PM BRF8 tripped off on a power amplifier air flow indication. Operations contacted an RF expert. He adjusted BRF19 to compensate for BRF8 being off.

At 4:42 PM, the Booster East Anode Power Supply (EAPS) tripped off due to a problem with BRF1.

At 4:44 PM, the NuMI horn tripped off. Operations tried to contact experts, but couldn’t get a hold of anyone. Operators investigated further and discovered that the horn had slipped into a conditioning mode. Operators reset the alarm and had the horn on by 5:30 PM.

At 5:54 PM, a RF expert reported that the modulator deck power had failed for BRF1. Since replacing this modulator would hold off beam, RF techs will access the Booster to repair BRF8 during the BRF1 repair.

RF experts reported at 6:34 PM that they had replaced the air filters for BRF8.

The Booster power supply work ended. Booster came back on and Pbar, NuMI, and MiniBooNE resumed taking beam at 7:30 PM.

The Recycler began preparing to stash at 8:13 PM. The stash ended and Pbar resumed stacking at 8:43 PM.

At 10:44 PM, the 400 MeV area Debuncher tripped off. It reset.

The Recycler began preparing to stash at 11:59 PM.

Friday July 28

The midnight shift began with Operations monitoring store 4859 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The stash ended at 12:32 AM.

At 1:02 AM, the NuMI horn tripped off on a PE1-A power supply water bit for the first of five times in about seven hours. It reset each time.

The Recycler began preparing to stash at 4:53 AM. The stash ended and Pbar resumed stacking around 5:30 AM.

The Plans for Friday and the Weekend

The plans for today are to continue monitoring store 4859. At 11 AM, the Run Coordinator will allow the TeV sector A1 pond strainer work to start as well as switching the feeders back to their normal configurations. Feeder 49 had a bad splice and was repaired yesterday. CDF and D0 will turn off during this work. The plans for the weekend are to stack, stash, and store.

Machine Reports

Linac

Current Status

  • Running on H- Source
    • Utilizing turbo pump and partial capacity of ion pump.
    • A new ion pump should be ready in the next couple of weeks at which point we will switch to the I- System.
  • Source is in on but in standby
  • Beam Current ~ 33 mA
  • Loss Sum Average ~ 13

Downtime Statistics

  • 95.6% Availability
  • 7.45 Hours with 25 Occurrences

Major Downtime

  • 1.28 Hours – LRF2 Modulator Sparking (Bad Resistor found)
  • 3.58 Hours – LRF4 Modulator (Power Outage Restart)
  • 1.07 Hours – LRF5 Driver Crowbar

Booster

Last Week

  • Stacking: 10-11 Turns, ~3.8 to 4.2E12 per batch
  • MiniBooNE: 11-12 Turns
  • NuMI: Starting up 8-9 Turns

Studies

  • Worked on higher intensity (12 Turn) ramps
  • Worked on Dampers

Downtime: 4 hours

  • RF stations 1 & 8 and misc. RF trips
  • Power glitch recovery

Next Week

Main Injector

MI operation

  • Shot to the TeV
    • Pbar coalescing eff. 84 - 88%
  • Shot to the RR
    • Efficiency 90 ~ 100%
  • Stacking
    • Intensity @injection ~ 7.4E12
    • Intensity @ target ~6.9E12 (Acc. eff. ~ 93%)
    • Tune up for mixed mode $23 cycle.

BPM upgrade

  • MI60 North electronics installation - complete
  • MI10 & MI50 electronics installation - next week

BLM upgrade

  • Still investigating the noise issues

MI studies

Antiproton Source

Last Week

  • Repaired Debuncher cooling horizontal & momentum band one kicker
  • Ran cables for PLC external power supply (4 of 6 PLCs)
  • Resurrected 120 GeV Overthruster (P1 BPMs masked)
  • 120 GeV Beamline optics measurements
  • AP3 trajectory tweaks
  • Corrected ARF4-MIRF energy mismatch for recycler transfers

Path to 20 mA/hr

  • Beam on target
    • 7.1E12 ® 7.5E12 => 5%
  • Beamline auto orbit correction (a.k.a. Overthruster)
    • 120 GeV BPM DAQ problems not resolved yet
    • Run 120 GeV Overthruster with P1 BPMs masked >> DONE
    • AP2 correction – requires stable 120 GeV orbit
    • This will reduce the large variation in stacking rate
      • Best stacking rate will not change
      • Average stacking rate will increase
  • Debuncher cooling kicker repair >> DONE
  • Fix stacktail cooling
    • Will allow the Accumulator to stack the increased flux when all of the above fixes are complete

Plans

  • Continue trouble-shooting 120 GeV beamline BPMs
    • Overthruster in use with P2 and AP1 BPMs
    • When BPMs fixed, re-establish full 120 GeV and AP2 auto orbit correction
  • Stacktail adjustments to increase Ed
  • Continue AP3 beamline adjustments

Recycler

Machine Status

  • Larger stashes
  • DCCT does not work above ~150E10
  • Requires 1 hour tunnel time
  • Use calibrated longitudinal Schottky power
  • Flying wire computer problem; fixed
  • Comparison of stores 4851, 4857 and 4859

Current Criteria for Tevatron Shots

  • What do we ask the operators to do?
    • Less than 5 π-mm-mr transverse emittances (Schottky)
    • Between 55 and 60 eVs independent of intensity
  • Are these numbers reasonable?

Studies

  • Pelletron studies: solenoids, drag rate (Ecool)
  • Pelletron energy alignment study
  • Flying wire sectional beam profile - preview
  • Transverse distribution vs. cooling -preview
  • Shot parameters – on-going

Tevatron

Store Summary

Store

Avg Init. Lum (E30)

CDF Deliv’d

Lum

(nb -1)

Time (hr)

How did it end?

Comments

4859

168.3

   

Ongoing

 

4857

142.3

4640

26.5

Intentional

Bunch lengthening; Tune problem; Abort gap loss problem; Bunch clipped by TEL1; After power glitch recovery;

4851

167.7

7761

40.8

Intentional

2nd highest luminosity; 1st avg. integral luminosity;

4847

148.4

6377

33.4

Intentional

Orbit stabilization program hang up;

4845

102.1

3571

24.9

Intentional

 

Machine Maintenance and Studies

  • New decoupling program in use
  • STRUCT system for closure
  • IPM testing
  • Flying wire checks
  • EOS beam tickler test
  • TEL2 modulator water flow meter replaced
  • C49 bump to reduce injection beam loss

Problems

  • E2 temperature high;
  • C:S3E3A feed down regulator replaced;
  • New version CLIB cause control system problem;
  • TRF7 modulator replaced;
  • E2A QBS controller replaced;
  • Power glitch:
    • Turbo pump checks;
    • F17 kicker 465 card void => F26U quench
  • D0 luminosity counter
  • Bunch length blowup; Abort gap loss grow;

Study Requests

  • Proton bumper timing adjustment:
    • Between stores (1 hr)
  • A0 beam abort position adjustment: EOS (1 hr)
  • TEL2 orbit and timing tuning, modulator test
    • Available proton only times
  • Timing E17 vertical kicker (Shot setup 30 min)
  • Recommissioning C100 (proton only 1~2 hr)
  • Helix orbit Measurement (Shot setup time 30 min)
  • Helix orbit tuning for HEP and ramp
  • Digital tune meter (EOS 30 min )
  • Vacuum measurement via beam lifetime (proton only 2~4 hr)

NuMI

Beam Returns

  • First beam Wednesday PM
  • Startup problems
    • One permit system C200 reported wrong data
    • TCLK fanout for power supply ramps failed badly
      • Made supplies ramp irregularly
      • Probably the cause of filter failure on Lambertson supply
    • BSCLK signal to BPMs bad. Swapped out
    • Beam position problems

Thanks for Help

  • Heartiest thanks from all of NuMI/MINOS to Mechanical Support Department for great job of cleaning up clogged water lines
    • Hundreds of tiny resin beads removed safely from an inaccessible location


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/28/2006    email Fermilab